# Interacting Multiple Model-based Joint Homography Matrix and Multiple Object State Estimation

### v1

@@ Abstract
A novel MOT algorithm, IMM Joint Homography State Estimation (IMM-JHSE), is proposed. By jointly modelling the camera projection matrix as part of track state vectors, IMM-JHSE removes the explicit influence of camera motion compensation techniques on predicted track position states, which was prevalent in previous approaches. Expanding upon this, static and dynamic camera motion models are combined through the use of an IMM filter. A simple bounding box motion model is used to predict bounding box positions to incorporate image plane information. In addition to applying an IMM to camera motion, a non-standard IMM approach is applied where bounding-box-based BIoU scores are mixed with ground-plane-based Mahalanobis distances in an IMM-like fashion to perform association only. Finally, IMM-JHSE makes use of dynamic process and measurement noise estimation techniques. IMM-JHSE improves upon related techniques on the DanceTrack and KITTI-car datasets, increasing HOTA by 2.64 and 2.11, respectively, while offering competitive performance on the MOT17, MOT20 and KITTI-pedestrian datasets.
@@ 1 Introduction
The increasing reliability of object detectors [1, 2, 3, 4, 5, 6, 7] has encouraged research on the topic of tracking objects in the image space, i.e. where objects are tracked in pixel coordinates. Most recent methods have focused on using bounding box information with or without appearance embeddings [8, 9, 10, 11, 12]. Yet, it is possible in some cases to obtain an estimate of the location of a bounding box within the ground plane of the imaged 3D space [13]. The ground plane is a coordinate system that represents a 2D approximation of the earth’s surface in front of the camera, as viewed from above. Leveraging such information may improve association performance. Occlusion and camera motion effects may be mitigated if the target dynamics within the ground plane are reliably estimated and the homographic projection of the ground plane to the image plane is decoupled from the target’s motion. This is illustrated in Figure 1, which depicts two tracks that undergo heavy occlusion in the image plane at time $t$. Although the red track is occluded by the blue track, their locations on the ground plane are still easily separable. Furthermore, camera motion does not influence target association, i.e. additional camera motion compensation is not required if an accurate projection matrix (homography) can be determined at any time step.
@@
    In addition to applying the IMM to camera motion as mentioned in the point directly above, the proposed method dynamically mixes ground-plane- and image-plane-based association scores, i.e. Mahalanobis distance and buffered intersection over union (BIoU) respectively, for the purpose of association only.
@@
The remainder of the paper is structured as follows. Section 2 examines previous work on the topic of image-based tracking. Section 3 provides more detail on motion models and camera motion compensation in the literature. Section 4 gives a detailed overview of the proposed method. Section 5 provides the experimental setup and results of tests on validation data. Section 6 reports the results on the MOT17, MOT20, DanceTrack and KITTI test datasets. Finally, Section 7 concludes the findings of this paper and suggests topics of focus for future work.
@@ 2 Previous Work
This section provides an overview of previous approaches to the MOT problem in order to provide context for the contributions presented in this paper.
@@ 2.1 Image-based Tracking
Tracking the motion of objects from a video has been of interest as early as 1998 when Isard and Blake [19] developed the condensation algorithm for tracking objects represented by a curve or sets of curves with learned motion models. Their method propagates conditional density distributions similar to the approach of particle filtering. However, tracking multiple objects simultaneously with their method may become infeasible, especially for complex multi-modal posterior densities where large samples or particles may be required to obtain accurate estimates. Besides, multi-object tracking requires solving the association problem – deciding which observations belong to which tracked object – which becomes increasingly difficult when the tracked objects are near one another, partially or fully occluding each other. Multiple object tracking (MOT) methods presented in the literature focus more on the association problem than on detection or individual target tracking.
@@ 2.2 Tracking Involving Homography
Various methods have used homography between multiple camera views to increase the robustness of the tracking of sports players. Since a particular player may be occluded in one camera view but not in another, using multiple cameras increases the robustness of a player tracking system to player occlusions. Examples of methods which make use of inter-camera homographies may be found in [37, 38, 39, 40, 41, 42]. However, few approaches consider the homography between a single camera view and the playing field. Notably, Hayet et al. [43] incrementally update the homography between the camera view and an imaged soccer field and explicitly use it to track player locations on the playing field with a Kalman filter. Recently, Maglo et al. [44] compute the homography to display player positions, but the tracking is performed without the aid of the homography. Instead, they rely heavily on appearance features extracted by a model fine-tuned offline after tracklet generation (on the corresponding bounding boxes). As a result, their method is not suitable for online operation. Most recently, Yi et al. [14] manually annotated the homographic projections between the image and ground planes of various datasets. They then perform tracking and association in the ground plane. However, as will be shown in detail in Section 4, camera motion explicitly influences target position in their method – which this paper regards as a modelling error.
@@ 3 Background: Camera Motion Compensation and Motion Models
$x^{I}_{t}=x^{I}_{t-1}+\dot{x}^{I}_{t-1}\Delta t,$ (1)
where $x^{I}_{t}$ represents the centre x- or y-coordinate, width or height (in pixels) of the target’s bounding box (these are jointly estimated and form part of the same state vector) at time $t$. The superscript $I$ indicates that the state vector $x^{I}_{t}$ is represented in the image plane. The individual state components will subsequently be denoted by $x^{I,x}_{t}$, $x^{I,y}_{t}$, $x^{I,w}_{t}$, $x^{I,h}_{t}$, respectively. The velocity of a state component is denoted by $\dot{x}$, and $\Delta t$ represents the sampling period — in this case, it is equal to the reciprocal of the video frame rate. The additive noise term has been omitted for brevity.
The current state is conditioned on the previous state, with measurements in the form of detected bounding boxes being directly related to the state itself. This model is graphically represented in Figure 2,
@@
Figure 2: The graphical model that is usually employed in modern MOT approaches. $x^{I}_{t}$ represents the state of a bounding box element (x/y position, width or height), and $y^{I}_{t}$ denotes the corresponding measurement, i.e. an element of the detected bounding box.
@@
$\mathbf{A}_{t}=\left[\begin{array}[]{c|c}\mathbf{R}^{A}_{t}\in\mathcal{R}^{2\times 2}&\mathbf{t}_{t}\in\mathcal{R}^{2\times 1}\\ \mathbf{0}^{1\times 2}&1\end{array}\right]$ (2)
@@
$\begin{bmatrix}x^{I,x^{\prime}}_{t}\\ x^{I,y^{\prime}}_{t}\\ 1\end{bmatrix}=\mathbf{A}_{t}\begin{bmatrix}x^{I,x}_{t}\\ x^{I,y}_{t}\\ 1\end{bmatrix}$ (3)
@@
$\begin{bmatrix}x^{I,w^{\prime}}_{t}\\ x^{I,h^{\prime}}_{t}\\ \end{bmatrix}=\mathbf{R}^{A}_{t}\begin{bmatrix}x^{I,w}_{t}\\ x^{I,h}_{t}\\ \end{bmatrix},$ (4)
where the prime superscript denotes the camera-motion-compensated state component. $\mathbf{A}_{t}$ is estimated with the camera calibration module of the OpenCV [45] library. Most previous approaches make use of optical flow features and the RANSAC [46] algorithm, although [14] makes use of the enhanced correlation coefficient (ECC) optimisation method.
@@
$x^{G}_{t}=x^{G}_{t-1}+\dot{x}^{G}_{t-1}\Delta t+w^{G},$ (5)
where $x^{G}_{t}$ represents the target’s x- or y-coordinate, and the superscript $G$ indicates that the state is represented in the ground plane. These components will subsequently be denoted by $x^{G,x}_{t}$ and $x^{G,y}_{t}$, respectively. The additive noise term $w^{G}$ is sampled from a zero-mean multivariate normal distribution with covariance $\mathbf{Q}^{G,x}$ (i.e. the process noise covariance matrix) determined by[14]:
$\mathbf{Q}^{G,x}=\mathbf{G}\operatorname{diag}\left(\sigma_{x},\sigma_{y}\right)\mathbf{G}^{T},$
@@
$\mathbf{G}=\left[\begin{array}[]{ll}\frac{\Delta t^{2}}{2}&0\\ \Delta t&0\\ 0&\frac{\Delta t^{2}}{2}\\ 0&\Delta t\end{array}\right],$
and $\sigma_{x}$ and $\sigma_{y}$ are process noise compensation factors along the x- and y-axes, respectively. The superscript $x$ in $\mathbf{Q}^{G,x}$ differentiates this noise term from those associated with the homography components in (8), (9).
@@
$\begin{bmatrix}x^{I,x}_{t}\\ x^{I,y}_{t}+0.5\times x^{I,h}_{t}\\ 1\end{bmatrix}=\operatorname{norm}\left(\mathbf{H}^{L}\begin{bmatrix}x^{G,x}_{t}\\ x^{G,y}_{t}\\ 1\end{bmatrix}\right),$ (6)
@@
$\mathbf{x}^{G^{\prime}}_{t}=\operatorname{norm}\left(\left(\mathbf{H}^{L}\right)^{-1}\mathbf{A}_{t}\operatorname{norm}\left(\mathbf{H}^{L}\mathbf{x}^{G}_{t}\right)\right).$ (7)
One might interpret (7) to imply that the target position in the ground plane is in some way dependent on camera motion, which is counter-intuitive to the knowledge that a target’s movement does not depend on camera motion. Target and camera motion models are treated independently in the proposed solution to address this.
@@ 4 The Proposed Solution
Let $\mathbf{H}^{G}_{t}$ represent the projection matrix as it evolves over time. Its evolution may be described by [13]
$\mathbf{H}^{G}_{t}=\mathbf{A}_{t}\mathbf{H}^{G}_{t-1}+\mathbf{\tilde{Q}_{t}^{G,H}},$ (8)
with $\mathbf{H}^{G}_{0}=\mathbf{H}^{L}$, and $\mathbf{\tilde{Q}_{t}^{G,H}}$ is an additive noise term which is estimated as described in Section 4.3. The superscript $\mathbf{H}$ differentiates this noise term from that of the ground plane position components, and the tilde differentiates it from the static motion model introduced in (9). Equation 8 describes how the homographic projection changes as a function of the previous projection and the estimated camera motion. Importantly, this is independent of and does not influence target motion.
A graphical model which combines Equations 5 and 8 with the measurement model in (6) is shown in Figure 3. Here, $x^{I}_{t}$ represents the bottom-centre bounding box position of the target detection. According to Bayesian network theory, the target and camera projection dynamics are independent while the target bounding box is unobserved, but they become dependent when conditioned on $x_{t}^{I}$ [48]. Thus, the temporal evolution of the ground plane and homography states are independent but are related through a common measurement in the image plane. This model may allow for a more accurate estimation of a target’s ground plane motion since its ground plane position prediction is not directly influenced by camera motion — only the position of its projection in the image plane.
$\mathbf{x}^{G}_{0}$$\mathbf{H}^{G}_{0}$$\mathbf{x}^{I}_{0}$cdots $\cdots$ $\mathbf{x}^{G}_{t}$$\mathbf{H}^{G}_{t}$$\mathbf{x}^{I}_{t}$
@@
The authors of [14] note that in scenes with relatively low camera motion, their method performs better without the camera motion compensation step in (7). This may be because the estimated affine matrix 2 is inaccurate, and noise is not considered in (7). To remedy this, the method proposed in this paper includes an additive noise term in (8) (see Sections 4.1 and 4.3). Additionally, to better model the evolution of $\mathbf{H}^{G}_{t}$ when there is no camera motion, the following motion model is included:
$\mathbf{H}^{G}_{t}=\mathbf{H}^{G}_{t-1}+\mathbf{\bar{Q}_{t}^{G,H}},$ (9)
where $\mathbf{\bar{Q}_{t}^{G,H}}$ is an additive noise term, where the bar differentiates it from that in (8). For each track, two filters are run in parallel. One of these makes use of (8), and the other makes use of (9). The states of these filters are fused in an IMM filter.
Furthermore, it has been shown that making use of buffered intersection over union (BIoU) during association can increase performance in scenes where either the camera or the tracks are highly dynamic [12]. Besides, there may be times when the tracks leave the ground plane (for example, jumps or flips in the DanceTrack dataset [15]), causing their measured ground plane positions to fluctuate wildly. During these times, it may be better to perform association with an image-based association measure such as the BIoU. Even if tracks do not leave the ground plane, they may be occluded, which could cause partial detections that manifest similar behaviour to when tracks leave the ground plane. Nevertheless, considering bounding box dimensions during association increases performance, as shown in Section 6. Therefore, the proposed method makes use of the simple bounding box motion model proposed in [12]. To perform association, the ground plane and image-based association scores are taken into account proportionally according to the predicted model probability [49, 50, 51]:
@@
Finally, the proposed method makes use of dynamic measurement and process noise estimation, which significantly improves performance (Section 4.3).

### v2

# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

@@ Abstract
A novel online MOT algorithm, IMM Joint Homography State Estimation (IMM-JHSE), is proposed. IMM-JHSE uses an initial homography estimate as the only additional 3D information, whereas other 3D MOT methods use regular 3D measurements. By jointly modelling the homography matrix and its dynamics as part of track state vectors, IMM-JHSE removes the explicit influence of camera motion compensation techniques on predicted track position states, which was prevalent in previous approaches. Expanding upon this, static and dynamic camera motion models are combined using an IMM filter. A simple bounding box motion model is used to predict bounding box positions to incorporate image plane information. In addition to applying an IMM to camera motion, a non-standard IMM approach is applied where bounding-box-based BIoU scores are mixed with ground-plane-based Mahalanobis distances in an IMM-like fashion to perform association only, making IMM-JHSE robust to motion away from the ground plane. Finally, IMM-JHSE makes use of dynamic process and measurement noise estimation techniques. IMM-JHSE improves upon related techniques, including UCMCTrack, OC-SORT, C-BIoU and ByteTrack on the DanceTrack and KITTI-car datasets, increasing HOTA by 2.64 and 2.11, respectively, while offering competitive performance on the MOT17, MOT20 and KITTI-pedestrian datasets. Using publicly available detections, IMM-JHSE outperforms almost all other 2D MOT methods and is outperformed only by 3D MOT methods—some of which are offline—on the KITTI-car dataset. Compared to tracking-by-attention methods, IMM-JHSE shows remarkably similar performance on the DanceTrack dataset and outperforms them on the MOT17 dataset. The code is publicly available: https://github.com/Paulkie99/imm-jhse.
@@ 1 Introduction
The increasing reliability of object detectors [1, 2, 3, 4, 5, 6, 7] has encouraged research on the topic of tracking objects in the image space, i.e. where objects are tracked in pixel coordinates. Most recent methods have focused on using bounding box information with or without appearance embeddings [8, 9, 10, 11, 12]. Yet, it is possible in some cases to obtain an estimate of the location of a bounding box within the ground plane of the imaged 3D space [13]. The ground plane is a coordinate system representing a 2D approximation of the earth’s surface in front of the camera, as viewed from above. Leveraging such information may improve association performance. Occlusion and camera motion effects may be mitigated if the target dynamics within the ground plane are reliably estimated and the homographic projection of the ground plane to the image plane is decoupled from the target’s motion. This is illustrated in Figure 1, which depicts two tracks that undergo heavy occlusion in the image plane at time $t$. Although the red track is occluded by the blue track, their locations on the ground plane are still easily separable. Furthermore, camera motion does not influence target association, i.e. additional camera motion compensation is not required if an accurate projection matrix (homography) can be determined at any time step.
In addition, many 3D MOT methods use regular 3D measurements [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], assume the internal camera parameters are known [29, 30] or attempt to estimate 3D coordinates based on 2D images [31]. Since IMM-JHSE requires a single initial homography estimate (which only provides a mapping to the 2D world plane), it is considered distinct from these approaches.
@@
    In addition to applying the IMM to camera motion as mentioned in the point directly above, the proposed method dynamically mixes ground-plane- and image-plane-based association scores, i.e. Mahalanobis distance and buffered intersection over union (BIoU), respectively, for association only.
@@
The remainder of the paper is structured as follows. Section 2 examines previous work on image-based tracking. Section 3 provides more detail on motion models and camera motion compensation in the literature. Section 4 gives a detailed overview of the proposed method. Section 5 provides the experimental setup and results of tests on validation data. Section 6 reports the results on the MOT17, MOT20, DanceTrack and KITTI test datasets. Finally, Section 7 concludes the findings of this paper and suggests topics of focus for future work.
@@ 2 Previous Work
This section provides an overview of previous approaches to the MOT problem to provide context for the contributions presented in this paper. Before reviewing the most relevant literature in Section 2.3 and Section 2.4, an overview of 3D MOT methods is provided in Section 2.1 to differentiate 3D MOT methods from the proposed solution. This is followed by a brief overview of tracking-by-attention methods in Section 2.2.
2.1 3D MOT
IMM-JHSE uses only a single initial homography estimate as in [32]. It is important to note that the homography transformation only provides a plane-to-plane mapping. This differentiates it from fully-3D MOT methods which use regular 3D measurements obtained with LiDAR sensors (point clouds) [14, 15, 16, 17, 18, 19], RGB-D cameras [37, 38], 3D detections [20, 21, 22, 23, 24, 25, 26], odometry [28] and stereo or binocular cameras [27, 39, 40].
In addition, the methods in [29, 30] use a Reversible Jump Markov Chain Monte Carlo particle filter with various feature representations, including skin colour, depth images, facial detection [29] and explicit occlusion reasoning [30]. Furthermore, these methods consider different camera motion models than the one proposed here (which explicitly considers camera position and rotation) and assume that internal camera parameters are known. Finally, [30] further makes use of odometry information.
Furthermore, the method in [31] attempts to perform 3D tracking without any of these measurements. However, the authors take an iterative approach to camera motion estimation. In contrast, IMM-JHSE integrates camera motion and its dynamics into the state vector to estimate it jointly with track states. They also rely on detecting vanishing points and lines parallel to the horizon. In addition, the multi-mode filter they employ only considers the target state and needs to be reinitialised with previous measurements when a mode switch occurs. Finally, multiple hypothesis tracking (MHT) performs association, whereas IMM-JHSE uses a linear assignment algorithm.
It could be argued that homography information is akin to 3D measurements. However, it should be emphasised that only one initial homography estimate is required for IMM-JHSE as opposed to the regular measurements exploited by 3D MOT methods. In addition, the homography is only a plane-to-plane mapping. As such, the authors of this paper still consider IMM-JHSE a 2D MOT method since two 2D planes are considered.
2.2 Tracking by attention
Instead of the pre-defined anchor boxes used in several popular object detectors, Carion et al. [7] decode object positions from a set of learned position embeddings using a transformer decoder. Meinhardt et al. [41] extend this approach to MOT by using previous detections as queries in subsequent frames, resulting in the Trackformer architecture. Whereas Trackformer works on a frame-by-frame basis, Global Tracking Transformers, as proposed by Zhou et al. [42], encode detections in a sliding temporal window of frames. A transformer decoder then associates new detections with these encodings. Having access to an entire measurement sequence, the deep data association module proposed by Pinto et al. [43] eschews the transformer decoder and relies solely on an encoder to associate detections within the given sequence with one another. End-to-end attention-based methods have recently seen success on MOT leaderboards [44, 45, 46]. However, these methods are computationally expensive, especially considering they often train their own detectors in addition to training the attention-based mechanism in an end-to-end approach.
@@ 2.3 Image-based Tracking
Tracking the motion of objects from a video has been of interest as early as 1998 when Isard and Blake [47] developed the condensation algorithm for tracking objects represented by a curve or sets of curves with learned motion models. Their method propagates conditional density distributions similar to the approach of particle filtering. However, tracking multiple objects simultaneously with their method may become infeasible, especially for complex multi-modal posterior densities where large samples or particles may be required to obtain accurate estimates. Besides, multi-object tracking requires solving the association problem – deciding which observations belong to which tracked object – which becomes increasingly difficult when the tracked objects are near one another, partially or fully occluding each other. Multiple object tracking (MOT) methods in the literature focus more on the association problem than on detection or individual target tracking.
@@ 2.4 Tracking Involving Homography
Various methods have used homography between multiple camera views to increase the robustness of the tracking of sports players. Since a particular player may be occluded in one camera view but not in another, using multiple cameras increases the robustness of a player tracking system to player occlusions. Examples of methods which make use of inter-camera homographies may be found in [65, 66, 67, 68, 69, 70]. However, few approaches consider the homography between a single camera view and the playing field. Notably, Hayet et al. [71] incrementally update the homography between the camera view and an imaged soccer field and explicitly use it to track player locations on the playing field with a Kalman filter. Recently, Maglo et al. [72] compute the homography to display player positions, but the tracking is performed without the aid of the homography. Instead, they rely heavily on appearance features extracted by a model fine-tuned offline after tracklet generation (on the corresponding bounding boxes). As a result, their method is not suitable for online operation. Most recently, Yi et al. [32] manually annotated the homographic projections between the image and ground planes of various datasets. They then perform tracking and association in the ground plane. However, as will be shown in Section 4, camera motion explicitly influences target position in their method – which this paper regards as a modelling error.
@@ 3 Background: Camera Motion Compensation and Motion Models
$\mathbf{x}^{I}_{t}=\mathbf{x}^{I}_{t-1}+\dot{\mathbf{x}}^{I}_{t-1}\Delta t,$ (1)
where $\mathbf{x}^{I}_{t}=\begin{bmatrix}x^{I}_{t}&y^{I}_{t}&w^{I}_{t}&h^{I}_{t}\end{bmatrix}^{\top}$ represents the centre x- and y-coordinate, width and height (in pixels) of the target’s bounding box at time $t$, respectively. The superscript $I$ indicates that the state vector $\mathbf{x}^{I}_{t}$ is represented in the image plane. The vector encompassing the velocity of each state component is denoted by $\dot{\mathbf{x}}$, and $\Delta t$ represents the sampling period — in this case, it is equal to the reciprocal of the video frame rate. The additive noise term has been omitted for brevity.
The current state is conditioned on the previous state, with measurements in the form of detected bounding boxes directly related to it. This model is graphically represented in Figure 2,
@@
Figure 2: The graphical model usually employed in modern MOT approaches. $\mathbf{x}^{I}_{t}$ represents the state of a bounding box element (x/y position, width or height), and $\mathbf{y}^{I}_{t}$ denotes the corresponding measurement, i.e. an element of the detected bounding box.
@@
$\mathbf{\hat{A}}_{t}=\left[\begin{array}[]{c|c}\mathbf{R}^{A}_{t}\in\mathcal{R}^{2\times 2}&\mathbf{t}_{t}\in\mathcal{R}^{2\times 1}\\ \mathbf{0}^{1\times 2}&1\end{array}\right]$ (2)
@@
$\begin{bmatrix}x^{I^{\prime}}_{t}\\ y^{I^{\prime}}_{t}\\ 1\end{bmatrix}=\mathbf{\hat{A}}_{t}\begin{bmatrix}x^{I}_{t}\\ y^{I}_{t}\\ 1\end{bmatrix}$ (3)
@@
$\begin{bmatrix}w^{I^{\prime}}_{t}\\ h^{I^{\prime}}_{t}\\ \end{bmatrix}=\mathbf{R}^{A}_{t}\begin{bmatrix}w^{I}_{t}\\ h^{I}_{t}\\ \end{bmatrix},$ (4)
where the prime superscript denotes the camera-motion-compensated state component. $\mathbf{\hat{A}}_{t}$ is estimated with the camera calibration module of the OpenCV [73] library as in [10]. Most previous approaches make use of optical flow features and the RANSAC [74] algorithm, although [32] makes use of the enhanced correlation coefficient (ECC) optimisation method.
@@
$\mathbf{x}^{W}_{t}=\mathbf{x}^{W}_{t-1}+\dot{\mathbf{x}}^{W}_{t-1}\Delta t+\mathbf{w}^{W},$ (5)
where $\mathbf{x}^{W}_{t}=\begin{bmatrix}x^{W}_{t}&y^{W}_{t}\end{bmatrix}^{\top}$ represents the target’s x- and y-coordinates, respectively. The superscript $W$ indicates that the state is represented in the ground (or world) plane. The additive noise term $\mathbf{w}^{W}$ is sampled from a zero-mean multivariate normal distribution with covariance $\mathbf{Q}^{W,\mathbf{x}}$ (i.e. the process noise covariance matrix) determined by[32]:
$\mathbf{Q}^{W,\mathbf{x}}=\mathbf{\Gamma}\operatorname{diag}\left(\sigma_{x},\sigma_{y}\right)\mathbf{\Gamma}^{T},$ (6)
@@
$\mathbf{\Gamma}=\left[\begin{array}[]{ll}\frac{\Delta t^{2}}{2}&0\\ \Delta t&0\\ 0&\frac{\Delta t^{2}}{2}\\ 0&\Delta t\end{array}\right],$ (7)
and $\sigma_{x}$ and $\sigma_{y}$ are process noise compensation factors along the x- and y-axes, respectively. The superscript $\mathbf{x}$ in $\mathbf{Q}^{W,\mathbf{x}}$ differentiates this noise term from those associated with the homography components in (10), (11).
@@
$\begin{bmatrix}x^{I}_{t}\\ x^{I}_{t}+0.5\times h^{I}_{t}\\ 1\end{bmatrix}=\operatorname{norm}\left(\mathbf{H}^{L}\begin{bmatrix}x^{W}_{t}\\ y^{W}_{t}\\ 1\end{bmatrix}\right),$ (8)
@@
$\mathbf{x}^{W^{\prime}}_{t}=\operatorname{norm}\left(\left(\mathbf{H}^{L}\right)^{-1}\mathbf{\hat{A}}_{t}\operatorname{norm}\left(\mathbf{H}^{L}\mathbf{x}^{W}_{t}\right)\right).$ (9)
One might interpret (9) to imply that the target position in the ground plane is in some way dependent on camera motion, which is counter-intuitive to the knowledge that a target’s movement does not depend on camera motion. Target and camera motion models are treated independently in the proposed solution.
@@ 4 The Proposed Solution
Let $\mathbf{H}^{W}_{t}$ represent the projection matrix as it evolves over time. Its evolution may be described by [13]
$\mathbf{H}^{W}_{t}=\mathbf{\mathbf{\hat{A}}}_{t}\mathbf{H}^{W}_{t-1}+\mathbf{\tilde{Q}}_{t}^{W,\mathbf{H}},$ (10)
with $\mathbf{H}^{W}_{0}=\mathbf{H}^{L}$, and $\mathbf{\tilde{Q}}_{t}^{W,\mathbf{H}}$ is an additive noise term which is estimated as described in Section 4.3. The superscript $\mathbf{H}$ differentiates this noise term from that of the ground plane position components, and the tilde differentiates it from the static motion model introduced in (11). Equation 10 describes how the homographic projection changes as a function of the previous projection and the estimated camera motion. Importantly, this is independent of and does not influence target motion. Note that the derivation of (10) in [13] assumes that ground plane points are static. While this is not the case where the target is a dynamic object, the motion model is applied nonetheless, especially since the process noise estimation procedure in Section 4.3 dynamically adjusts the expected error statistics.
A graphical model which combines Equation 5 and Equation 10 with the measurement model in (8) is shown in Figure 3. Here, $\mathbf{x}^{I}_{t}$ represents the bottom-centre bounding box position of the target detection. According to Bayesian network theory, the target and camera projection dynamics are independent while the target bounding box is unobserved, but they become dependent when conditioned on $\mathbf{x}_{t}^{I}$ [76]. Thus, the temporal evolution of the ground plane and homography states are independent but are related through a common measurement in the image plane. This model may allow for a more accurate estimation of a target’s ground plane motion since its ground plane position prediction is not directly influenced by camera motion — only the position of its projection in the image plane.
$\mathbf{x}^{W}_{0}$$\mathbf{H}^{W}_{0}$$\mathbf{x}^{I}_{0}$cdots $\cdots$ $\mathbf{x}^{W}_{t}$$\mathbf{H}^{W}_{t}$$\mathbf{x}^{I}_{t}$
@@
The authors of [32] note that in scenes with relatively low camera motion, their method performs better without the camera motion compensation step in (9). This may be because the estimated affine matrix (2) is inaccurate, and noise is not considered in (9). To remedy this, the method proposed in this paper includes an additive noise term in (10) (see Section 4.1 and Section 4.3). Additionally, to better model the evolution of $\mathbf{H}^{W}_{t}$ when there is no camera motion, the following motion model is included:
$\mathbf{H}^{W}_{t}=\mathbf{H}^{W}_{t-1}+\mathbf{\bar{Q}_{t}^{W,\mathbf{H}}},$ (11)
where $\mathbf{\bar{Q}_{t}^{W,\mathbf{H}}}$ is an additive noise term, and the bar differentiates it from that in (10). For each track, two filters are run in parallel. One of these makes use of (10), and the other makes use of (11). The states of these filters are fused in an IMM filter as described in A.
Furthermore, it has been shown that using buffered intersection over union (BIoU) during association can increase performance in scenes where the camera or the tracks are highly dynamic [12]. Besides, there may be times when the tracks leave the ground plane (for example, jumps or flips in the DanceTrack dataset [33]), causing their measured ground plane positions to fluctuate wildly. During these times, it may be better to perform association with an image-based association measure such as the BIoU. Even if tracks do not leave the ground plane, they may be occluded, which could cause partial detections that manifest similar behaviour to when tracks leave the ground plane. Nevertheless, considering bounding box dimensions during association increases performance, as shown in Section 6. Therefore, the proposed method uses the simple bounding box motion model proposed in [12]. To perform association, the ground plane and image-based association scores are taken into account proportionally according to the predicted model likelihood [77]:
@@
Finally, the proposed method uses dynamic measurement and process noise estimation, significantly improving performance (see Section 4.3).